# Theory-agnostic searches for non-gravitational modes in black hole ringdown

### v2

@@ A.2 The effect of precession on GW190521 and GW200129
There are well-founded arguments suggesting that GW190521 and GW200129 could originate from precessing binaries [81, 79]. As opposed to aligned-spin systems, precession breaks the symmetry between $m\leftrightarrow-m$ angular modes, resulting in $h_{lm}\neq(-1)^{l}h^{*}_{l-m}$. In the main text, we searched for a scalar mode using a $(lm)=(2,\pm 2)$ spin-aligned (non-precessing) waveform. Here, we extend the parameter estimation on GW190521 and GW200129 including precession, and varying $t_{\text{offset}}=[-2,0,2]\text{ms}$. This is achieved by allowing the amplitude $A_{22n}\neq A_{2-2n}$ and the phase $\phi_{22n}\neq-\phi_{2-2n}$, for both the gravitational and scalar mode. In Fig. 3, we show the corner plots for GW200129 with $t_{\text{offset}}=0$, for both the GR01 and GR0+S cases. Notice that, since only $(lm)=(2,\pm 2)$ modes are used, accounting for precession in this case does not significantly affect the values of the mass and the spin of the remnant compared to the aligned-spin case. However, including precession affects the amplitude of the fundamental mode, by increasing its uncertainty. We find qualitatively similar results for the other offset times and for GW190521.
A.3 Supplemental results
Here we present some results complementing those discussed in the main text.
@@ A.3 Supplemental results
Finally, in Fig. 6, we compare some representative posterior distributions obtained from the Bayesian inference on real data with forecasts using injections at higher SNR (equal to 100). We consider a ringdown model with an extra scalar mode (GR0+S, left panel) and with an extra vector mode (GR0+V, right panel). For the injection simulations, we inject the values corresponding to the maximum likelihood values of the real data. We notice that, at higher SNRs, the precision of the parameters’ distribution improves, and the signal is well reconstructed. This can be easily observed with the distributions of $\chi_{f}$, which in the case of the real data is spread, due to the effect of the gated Gaussian noise. Also in this case we note that the results for the scalar or vector case are very similar.

### v3

@@ VII.2 The effect of precession on GW190521 and GW200129
There are well-founded arguments suggesting that GW190521 and GW200129 could originate from precessing binaries [Gupta:2024gun, Siegel:2023lxl]. As opposed to aligned-spin systems, precession breaks the symmetry between $m\leftrightarrow-m$ angular modes, resulting in $h_{lm}\neq(-1)^{l}h^{*}_{l-m}$. In the main text, we searched for a scalar mode using a $(lm)=(2,\pm 2)$ spin-aligned (non-precessing) waveform. Here, we extend the parameter estimation on GW190521 and GW200129 including precession, and varying $t_{\text{offset}}=[-2,0,2]\text{ms}$. This is achieved by allowing the amplitude $A_{22n}\neq A_{2-2n}$ and the phase $\phi_{22n}\neq-\phi_{2-2n}$, for both the gravitational and scalar mode. In Fig. 3, we show the corner plots for GW200129 with $t_{\text{offset}}=0$, for both the GR1 and GR0+S cases. Notice that, since only $(lm)=(2,\pm 2)$ modes are used, accounting for precession in this case does not significantly affect the values of the mass and the spin of the remnant compared to the aligned-spin case. However, including precession affects the amplitude of the fundamental mode, by increasing its uncertainty. We find qualitatively similar results for the other offset times and for GW190521.
VII.3 Supplemental results
@@
Further, in Fig. 6, we compare some representative posterior distributions obtained from the Bayesian inference on real data with forecasts using injections at higher SNR (equal to 100). We consider a ringdown model with an extra scalar mode (GR0+S, left panel) and with an extra vector mode (GR0+V, right panel). For the injection simulations, we inject the values corresponding to the maximum likelihood values of the real data. We notice that, at higher SNRs, the precision of the parameters’ distribution improves, and the signal is well reconstructed. This can be easily observed with the distributions of $\chi_{f}$, which in the case of the real data is spread, due to the effect of the gated Gaussian noise. Also in this case, we note that the results for the scalar or vector case are very similar.
Finally, we have searched for multi-modality in the amplitude of the scalar mode. This is supported by the fact that, for some different choices of mass and spin, the fundamental scalar/vector mode can mimic the fundamental gravitational QNM. In general, this effect is not evident due to the limited range of the prior on the amplitude ratio $\hat{A}_{220}\in[0,10]$. We have thus performed a new run where we inject GR0 with GW150914-like parameters (but in our case, we have $SNR=83$), and recover with GR0+S, but with a broader prior on the amplitudes ($A_{220},\hat{A}_{R,220}\in[0,100]$), indeed finding bimodality. Including a secondary mode in the ringdown analysis, as we did in certain cases, clearly breaks this degeneracy.
@@
VII.4 The detectability and resolvability criteria
A BH no-hair test is designed to check the consistency of the mass and spin as observed from two different ringdown modes: the fundamental mode $(l=2,m=2,n=0)$, and a next-to-leading order mode with either $|l|=|m|\neq 2$ or $n\neq 0$. Ideally, one would have to infer to significant accuracy the tetrad of parameters $\left\{\omega_{220},\tau_{220},\omega_{lmn},\tau_{lmn}\right\}$. In practice, one just needs to build a triad $\left\{\omega_{220},\tau_{220},q_{lmn}\right\}$ with $q_{lmn}=\omega_{lmn}$ or $q_{lmn}=\tau_{lmn}$, and check the consistency on the inferred mass and the spin from the following tuples $\left\{(\omega_{220},\tau_{220}),(\omega_{220},q_{lmn}),(\tau_{220},q_{220})\right\}$ [Gossan:2011ha]. Whether we can reliably distinguish a secondary mode depends critically on its SNR. The threshold SNR can be defined in different ways as:
  1. 1.
    $\rho_{\text{det}}$: The ringdown SNR at which the amplitude of the mode is large enough compared to its statistical uncertainty $\sigma_{A_{R}}/A^{R}_{lmn}\sim 1$.
  2. 2.
    $\rho_{\text{res}}$: The ringdown SNR at which the spectrum $q_{lmn}$ subtracted to the spectrum $q_{220}$ of fundamental mode is comparable to its statistical uncertainty $\sigma_{q_{lmn}}/|q_{220}-q_{lmn}|\sim 1$ ,
where $\sigma_{A_{R}}$ and $\sigma_{q_{lmn}}$ are the $1-\sigma$} statistical errors on $A_{R}$ and $\sigma_{q_{lmn}}$ respectively [Forteza:2020hbw]. A mode is measured if conditions 1) and 2) are satisfied, which will occur for $\rho\gtrsim{\rm max}\left\{\rho_{\text{det}},\rho_{\text{res}}\right\}$ [Berti:2005ys, Forteza:2020hbw]. Using the Fisher Matrix approach, the set $\{\rho_{\text{det}},\rho_{\text{res}}\}$ can be estimated analytically, provided that the sensitivity curve in the ringdown regime is sufficiently flat for a given total mass. The statistical error $\sigma_{q_{lmn}}$ and the SNR of a signal $\rho$ are computed as,
$\sigma_{ij}^{2}\propto\bigg(\int{\frac{\frac{dh}{d\lambda_{i}}(\frac{dh}{d\lambda_{j}})^{*}}{S_{n}(f)}df}\bigg)^{-1}\quad,\quad\rho^{2}\propto\int\frac{|h|^{2}}{S_{n}(f)}df\,$, (13)
where $\lambda_{i,j}$ are each of the ringdown intrinsic parameters, denotes the complex conjugate and -1 the matrix inverse. Note that the dependence of the waveform on the of extrinsic parameters $f(\Theta)$ will be the same for both equations since one can replace $h\rightarrow h\cdot f(\Theta)$. The GW ringdown spans over a relatively short frequency range. For the scalar mode case, and for $M_{f}=62$, notice that the whole ringdown signal is approximately all contained within $\left[200,300\right]\text{Hz}$ at $a_{f}\sim 0.7$ (see Fig. 4). Therefore, one can assume an effective $S_{n}(f)~\sim S_{n}(f_{0})$ where $f_{0}$ is an appropriate frequency within the range described above. The last approximation allows us to define $f(\lambda_{ij})\equiv\rho_{(\text{det},\text{res})}\cdot\sigma_{ij}$, which only depends on the intrinsic parameters $\lambda_{ij}$, and allows us to also describe both the $\rho_{\text{det}}$ and $\rho_{\text{res}}$ independent from the intrinsic parameters as
$\rho_{\text{det}}=\frac{f(\lambda_{A^{R}_{lmn}})}{A^{R}_{lmn}}\quad,\quad\rho_{\text{res}}=\frac{f(\lambda_{q_{lmn}})}{|q_{220}-q_{lmn}|}\,.$ (14)
For scalar and vector modes, we obtain that $\rho_{\text{det}}\gg\rho_{\text{res}}$ for $A^{R}_{lmn}\lesssim 1$. Therefore, $\rho>\rho_{\rm det}$ is our determining criterion for observing them. The above equation also shows clearly the trend $\rho_{\text{det}}\sim(A^{R})^{-1}$ displayed in Fig. 2 of the main text. Similarly, a $M_{f}~\sim 10^{7}M_{\odot}$ binary will cover a short frequency range at the mHz, in which current estimates of the LISA sensitivity curve are rather flat [Colpi:2024xhw]. Finally, notice that both $\rho_{\text{det}}$ and $\rho_{\text{res}}$ scale with $f\propto\rho\cdot\sigma_{ij}$. For a binary system considering only the $(\ell,m)=(2,2)$ mode, non-precessing, and analyzed in the frequency domain, the GW strain can be expressed as $\tilde{h}=H_{22}(\vec{\Theta})h_{+}(\vec{\lambda})e^{i\phi_{22}(\vec{\lambda})}$ where $H_{22}$ is an amplitude factor that encodes the sky position, polarization angle, and distance, $h_{+}(\vec{\lambda})$ is the plus polarization component, $\phi_{22}$ is the phase, and $\vec{\lambda}$ represents the intrinsic parameters of the system [Babak:2012zx]. Therefore, it follows that the product $f\propto\rho\cdot\sigma_{ij}$ is independent of the system’s sky position, implying that both $\rho_{\text{det}}$ and $\rho_{\text{res}}$ are also independent of $\vec{\Theta}$.